# A geometric approximation of non-local interface and boundary conditions

### v1

@@ 3. Example
$\mathbf{\Gamma}_{s,\varepsilon}=\big{\{}\mathbf{x}=(x^{1}\dots,x^{n-1})\in\mathbb{R}^{n}:\ |x^{k}-\varepsilon s^{k}|<\varepsilon/2,\ k=1,\dots,n-1\big{\}}.$
@@
$\alpha_{s,t,\varepsilon}=(K(x_{s,t,\varepsilon},x_{t,s,\varepsilon}))^{1/(n-1)},\text{ where }x_{s,t,\varepsilon}=(\mathbf{x}_{s,t,\varepsilon},0)\in\Gamma$ (3.4)
@@
Evidently, $\ell_{\varepsilon}$ is a bijective map. Let the passage $T_{i,j,\varepsilon}$ with $i\in\mathbb{I}_{\varepsilon}$ and $j=\ell(i)$ be given by (2.6) with
@@
where $\Gamma_{s,\varepsilon}\coloneqq\{x\in\mathbb{R}^{n}:\ (x^{1},\dots,x^{n-1})\in\mathbf{\Gamma}_{s,\varepsilon},\ x^{n}=0\}$. By construction, the sets $\Gamma_{s,\varepsilon}\times\Gamma_{s,\varepsilon}$, $(s,t)\in\mathbb{Z}_{\varepsilon}\times\mathbb{Z}_{\varepsilon}$ are mutually disjoint, and we have
$(x_{s,t,\varepsilon},x_{t,s,\varepsilon})\in\Gamma_{s,\varepsilon}\times\Gamma_{t,\varepsilon},\ \forall(s,t)\in\mathbb{Z}_{\varepsilon}\times\mathbb{Z}_{\varepsilon}.$
Furthermore, due to (3.1), one has
@@ 4.1. Auxiliary estimates
$\leq C_{1}\|u_{\varepsilon}\|_{L^{2}(M_{\varepsilon})}\bigg{(}h_{\varepsilon}\sum_{i\in\mathbb{I}_{\varepsilon}}\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon}^{\pm})\bigg{)}^{1/2}\to 0\;\text{ as }\;\varepsilon\to 0.$ (4.45)
where the last conclusion follows from (4.29), (2.15) and the fact that the sets $\mathbf{D}_{i,\varepsilon}^{\pm}$ are pairwise disjoint belong to $\mathbf{\Gamma}$ (whence, their total volume is bounded). Similarly, taking into account (4.28), we get
@@
$\leq C_{1}\bigg{[}\sum_{i\in\mathbb{I}_{\varepsilon}}\frac{K_{i,\ell(i),\varepsilon}}{\operatorname{area}_{n-1}(B_{i,\varepsilon}^{+})}\|u_{\varepsilon}-\widetilde{u}^{+}_{\delta}\|^{2}_{L^{2}(B_{i,\varepsilon}^{+})}+\sum_{j\in\mathbb{I}_{\varepsilon}}\frac{K_{\ell^{-1}(j),j,\varepsilon}}{\operatorname{area}_{n-1}(B_{j,\varepsilon}^{-})}\|u_{\varepsilon}-\widetilde{u}^{-}_{\delta}\|^{2}_{L^{2}(B_{j,\varepsilon}^{-})}\bigg{]}^{1/2}$
@@
$\leq C_{1}\left(\sum_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{+})^{n-1}q_{i,\varepsilon}^{+}\|\nabla u_{\varepsilon}\|^{2}_{L^{2}(Y_{i,\varepsilon}^{+})}+\sum_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{-})^{n-1}q_{i,\varepsilon}^{-}\|\nabla u_{\varepsilon}\|^{2}_{L^{2}(Y_{i,\varepsilon}^{+})}\right)^{1/2}$
$\leq C_{2}\bigg{(}\big{(}\sup_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{+})^{n-1}q_{i,\varepsilon}^{+}+\sup_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{-})^{n-1}q_{i,\varepsilon}^{-}\big{)}\mathfrak{h}_{\varepsilon}[u_{\varepsilon},u_{\varepsilon}]\bigg{)}^{1/2}.$
@@
Furthermore, we also have $|v_{\varepsilon}(t)-v_{\varepsilon}(s)|^{2}\leq|t-s|\displaystyle\int_{0}^{\max\{s,t\}}|\partial_{t}v_{\varepsilon}(\tau)|^{2}\,\mathrm{d}\tau,$ whence in view of (2.27) and (4.73), we infer that
@@ 5. Non-local Robin boundary conditions
$\begin{cases}-\Delta u+u=f&\text{in }\Omega,\\ \displaystyle\frac{\partial u}{\partial x^{n}}+\int_{\Gamma}K(x,y)(u(x)-u(y)){\,\mathrm{d}s_{y}}=0&\text{on }\Gamma,\\[5.69054pt] \displaystyle\frac{\partial u}{\partial\nu}=0&\text{on }\partial\Omega\setminus\Gamma.\end{cases}$
@@
We left the proof of this claim to the reader, since it follows the proofs of Theorems 2.32.6 almost verbatim, with minor modifications only, such as adjustments in the construction of the test function $w_{\varepsilon}(x)$.
@@
Figure 5.2. The set $\mathbf{\Gamma}$. The big (respectively, small) cubes correspond to the sets $\mathbf{\Gamma}_{s,\varepsilon}$ (respectively, $\mathbf{\Gamma}_{s,t,\varepsilon}$). The dotted lines link the cubes $\mathbf{\Gamma}_{s,t,\varepsilon}$ and $\mathbf{\Gamma}_{t,s,\varepsilon}$ (i.e., those cubes that further will be linked by a passage). The black cubes corresponds to $\mathbf{\Gamma}_{s,t,\varepsilon}$ with $s=t$; within these cubes we have no ’holes’ $\mathbf{D}_{s,t,\varepsilon}$

### v2

@@ 3. Example
$\mathbf{\Gamma}_{s,\varepsilon}=\big{\{}\mathbf{x}=(x^{1}\dots,x^{n-1})\in\mathbb{R}^{n-1}:\ |x^{k}-\varepsilon s^{k}|<\varepsilon/2,\ k=1,\dots,n-1\big{\}}.$
@@
$\alpha_{s,t,\varepsilon}=(K(x_{s,t,\varepsilon},x_{t,s,\varepsilon}))^{1/(n-1)},\text{ where }x_{s,t,\varepsilon}\coloneqq(\mathbf{x}_{s,t,\varepsilon},0)\in\Gamma$ (3.4)
@@
Evidently, $\ell_{\varepsilon}$ is a bijective map. Let the passage $T_{i,j,\varepsilon}$ with $i\in\mathbb{I}_{\varepsilon}$ and $j=\ell_{\varepsilon}(i)$ be given by (2.6) with
@@
where $\Gamma_{s,\varepsilon}\coloneqq\{x\in\mathbb{R}^{n}:\ (x^{1},\dots,x^{n-1})\in\mathbf{\Gamma}_{s,\varepsilon},\ x^{n}=0\}$. By construction, we have
$(x_{s,t,\varepsilon},x_{t,s,\varepsilon})\in\Gamma_{s,\varepsilon}\times\Gamma_{t,\varepsilon},\ \forall(s,t)\in\mathbb{Z}_{\varepsilon}\times\mathbb{Z}_{\varepsilon},$
the sets $\{\Gamma_{s,\varepsilon}\times\Gamma_{s,\varepsilon},\ (s,t)\in\mathbb{Z}_{\varepsilon}\times\mathbb{Z}_{\varepsilon}\}$ are mutually disjoint, and one has
@@ 4.1. Auxiliary estimates
$\leq C_{1}\|u_{\varepsilon}\|_{L^{2}(M_{\varepsilon})}\bigg{(}h_{\varepsilon}\sum_{i\in\mathbb{I}_{\varepsilon}}\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon}^{+})\bigg{)}^{1/2}\to 0\;\text{ as }\;\varepsilon\to 0.$ (4.45)
where the last conclusion follows from (4.29), (2.15) and the fact that the sets $\mathbf{D}_{i,\varepsilon}^{+}$ are pairwise disjoint and belong to $\mathbf{\Gamma}$ (whence, their total volume is bounded). Similarly, taking into account (4.28), we get
@@
$\leq C_{1}\bigg{[}\sum_{i\in\mathbb{I}_{\varepsilon}}\frac{K_{i,\ell_{\varepsilon}(i),\varepsilon}}{\operatorname{area}_{n-1}(B_{i,\varepsilon}^{+})}\|u_{\varepsilon}-\widetilde{u}^{+}_{\delta}\|^{2}_{L^{2}(B_{i,\varepsilon}^{+})}+\sum_{j\in\mathbb{I}_{\varepsilon}}\frac{K_{\ell_{\varepsilon}^{-1}(j),j,\varepsilon}}{\operatorname{area}_{n-1}(B_{j,\varepsilon}^{-})}\|u_{\varepsilon}-\widetilde{u}^{-}_{\delta}\|^{2}_{L^{2}(B_{j,\varepsilon}^{-})}\bigg{]}^{1/2}$
@@
$\leq C_{1}\left(\sum_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{+})^{n-1}q_{i,\varepsilon}^{+}\|\nabla u_{\varepsilon}\|^{2}_{L^{2}(Y_{i,\varepsilon}^{+})}+\sum_{j\in\mathbb{I}_{\varepsilon}}(\varrho_{j,\varepsilon}^{-})^{n-1}q_{j,\varepsilon}^{-}\|\nabla u_{\varepsilon}\|^{2}_{L^{2}(Y_{j,\varepsilon}^{-})}\right)^{1/2}$
$\leq C_{2}\bigg{(}\big{(}\sup_{i\in\mathbb{I}_{\varepsilon}}(\varrho_{i,\varepsilon}^{+})^{n-1}q_{i,\varepsilon}^{+}+\sup_{j\in\mathbb{I}_{\varepsilon}}(\varrho_{j,\varepsilon}^{-})^{n-1}q_{j,\varepsilon}^{-}\big{)}\mathfrak{h}_{\varepsilon}[u_{\varepsilon},u_{\varepsilon}]\bigg{)}^{1/2}.$
@@
Furthermore, we also have $|v_{\varepsilon}(t)-v_{\varepsilon}(s)|^{2}=\displaystyle\left|\int_{s}^{t}\partial_{t}v_{\varepsilon}(\tau)\,\mathrm{d}\tau\right|^{2}\leq|t-s|\int_{0}^{\max\{s,t\}}|\partial_{t}v_{\varepsilon}(\tau)|^{2}\,\mathrm{d}\tau,$ whence in view of (2.29) and (4.73), we infer that
@@ 5. Non-local Robin boundary conditions
$\begin{cases}-\Delta u+u=f&\text{in }\Omega,\\ \displaystyle\frac{\partial u}{\partial x^{n}}=-\int_{\Gamma}K(x,y)(u(x)-u(y)){\,\mathrm{d}s_{y}}&\text{on }\Gamma,\\[5.69054pt] \displaystyle\frac{\partial u}{\partial\nu}=0&\text{on }\partial\Omega\setminus\Gamma.\end{cases}$
@@
We left the proof of this claim to the reader, since it follows the proofs of Theorems 2.3, 2.5, 2.6 almost verbatim, with minor modifications only, such as adjustments in the construction of the test function $w_{\varepsilon}(x)$.
@@
Figure 5.2. The set $\mathbf{\Gamma}$. The big (respectively, small) cubes correspond to the sets $\mathbf{\Gamma}_{s,\varepsilon}$ (respectively, $\mathbf{\Gamma}_{s,t,\varepsilon}$). The dotted lines link the cubes $\mathbf{\Gamma}_{s,t,\varepsilon}$ and $\mathbf{\Gamma}_{t,s,\varepsilon}$ (i.e., those cubes that further will be linked by a passage). The black cubes corresponds to $\mathbf{\Gamma}_{s,t,\varepsilon}$ with $s=t$; within these cubes we have no ‘holes’ $\mathbf{D}_{s,t,\varepsilon}$